# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v1

@@ I.4 Evaluation Method and Metrics
Experiment II: We quantify the model’s speaker preference, specifically its inclination to authenticate speakers using complex attack techniques, by tracking the proportional distribution of authentications between the two speakers presented.
Experiment III: We calculate the authentication pass rate.
Experiment IV: We compare authentication pass rates across different experimental groups.
@@ I.5 Additional Result
| MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | |
| Awake Assistant | 27 | 15 | N/A | 91 | 100 | 94 | 99 | 33 | 8 |
| Open The Door | 14 | 0 | N/A | 53 | 33 | 84 | 80 | 0 | 20 |
@@
Voice Cloning Spoofing. In Table 24, we analyze the experimental results of all open-source and closed-source models under two scenarios with different levels of text prompt flexibility. It can be observed that most models perform better in the “Open The Door” scenario than in the “Awake Assistant” scenario. with a significant decrease in the number of samples that failed to defend against voice cloning spoofing. This is particularly evident for Ultravox, Gemini-1.5 Pro, and Gemini-2.5 Pro. This indicates that even in Audio-based Large Language Models (ALLMs) where audio is the primary input, the accuracy of text prompts still plays a significant role. Furthermore, this has implications for the downstream applications of ALLMs: for scenarios involving security, authentication, etc., designing a strict and precise prompt may lead to a considerable improvement in model performance.
| MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | |
| Identity Verification Bypass | 76 | 58 | 74 | 5 | 4 | 3 | 5 | 2 | 0 |
| Identity Verification Bypass (Text) | x | x | x | x | 6 | 4 | 0 | 6 | 9 |
| Hybrid Spoofing(bg) | 80 | 30 | 92 | 50 | 13 | 19.8 | 15.5 | 10 | 7 |
| Hybrid Spoofing(no bg) | 80 | 41.4 | 90 | 46 | 10 | 28.4 | 10.4 | 9.1 | 7 |
| Speaker Preference | 81.8 | 59 | 39 | 72.4 | 96 | 80 | 81 | 100 | 100 |
@@
Identity Verification Bypass. From the Table 25, it can be observed that closed-source models are harder to deceive compared to open-source models. Among them, GPT-4o Audio performs the best, with a FAR (False Acceptance Rate) of only 2%. Among all closed-source models, SALMONN performs the worst, with a FAR as high as 76%. These results indicate that even without providing complete or explicit authentication information, voice models still have a high probability of passing identity verification, which poses a significant security risk.In the Table 25, we also investigated the FAR metric in pure text mode, which is labeled as the “Text” column. This represents using the corresponding text-based model of the audio model to perform inference on the text version of the identity verification bypass dataset. It can be observed that, in general, the FAR is higher in text mode compared to audio mode. This suggests that the additional paralinguistic information present in the speech modality, such as emotional cues or prosodic features, may contribute positively to the authentication performance of the model.
@@
Hybrid Spoofing. In this configuration, we simulate social engineering attacks combined with background audio that may occur in real authentication scenarios, aiming to study the impact of background sounds on the verification outcome. The experimental results show that the influence of added background audio on model performance does not follow an obvious pattern. For instance, Qwen2-Audio’s FAR decreases by 11.4%, whereas Gemini-1.5 Pro’s FAR increases instead.
Speaker Preference. To investigate how social engineering techniques affect speaker verification systems, we designed an experiment in which models were required to choose between two audio samples and identify the one more likely to belong to a legitimate user. The results show that SALMONN has a 39% tendency to select the audio from the genuine user for authentication. In contrast, the GPT-4o series of models consistently favors the audio from the attacker employing social engineering techniques, treating it as belonging to an authenticated user.
@@
With the rapid increase in parameter and data scales, text-only large language models (LLMs) have achieved groundbreaking progress in language understanding and generation, as exemplified by models such as GPT-4 and the Gemini series [1, 66]. Building on this, researchers explored cross-modal alignment by integrating visual information into unified representation spaces. This led to early models like CLIP [57] and Flamingo [3], and later, models such as GPT-4V and Gemini capable of processing high-resolution images and long contexts. Recently, ALLMs have further expanded the input modalities by incorporating temporal acoustic features (such as Mel-spectrograms, log-power spectra, or variable-length waveforms) for joint modeling with semantic tokens [86]. In contrast to the visual modality, audio signals exhibit high dynamic range and transient variations in both time and frequency domains. Consequently, most ALLMs adopt separate time-frequency encoders or discretizing acoustic tokenizers to capture rich attributes such as timbre, rhythm, and scene [27, 18]. Representative models include Qwen2-Audio with its pipeline-style natural language prompt pre-training [14], SALMONN with a unified "auditory-language-music" framework [65], and WavLLM with a dual-encoder plus Prompt-LoRA adaptation mechanism [27]. After cross-modal alignment, these models demonstrate strong capabilities in content and scene understanding, enabling applications such as spoken question answering, music style analysis, and environmental sound event retrieval. They also show great promise in medical diagnosis (e.g., detection of respiratory diseases, analysis of heart sounds), voice control for smart homes, and multimedia generation and editing [99, 59, 5].
@@
However, the multimodal nature of ALLMs also introduces new trustworthiness challenges. First, since the models are trained on large-scale acoustic-text paired corpora, they are prone to memorizing and leaking sensitive user speech information, and are therefore vulnerable to privacy attacks such as membership inference [67, 26]. Second, adversarial audio can exploit inaudible ultrasound or fine-grained perturbations to mislead ALLMs: early work such as DolphinAttack [96] and Vrifle [38] demonstrated covert manipulation of voice assistants via inaudible commands injected with ultrasonic carriers above 20 kHz [104, 36, 92]; recently, AdvWave systematically proposed gradient shattering repair and two-stage optimization, achieving over 40% jailbreak success rates on various ALLMs [30]. In addition, large-scale multimodal models are similarly susceptible to cross-modal instruction injection and protocol mismatching attacks, potentially leading to unauthorized content generation [43], privilege escalation [23], and even physical harm [45]. When integrated into voice-interface agentic systems, trustworthiness challenges are amplified and become paramount [40, 91]. To address these risks, the community has proposed a range of safety, security, and privacy mechanisms, including SafeEar, an empirical content privacy-preserving audio deepfake detection framework [37] and active detection with post-hoc rejection [35] differentially private pre-training, segment-wise gradient compression defenses. Nevertheless, in real-time voice scenarios, these approaches still face detection latency and robustness trade-offs, highlighting the urgent need for further research.
@@
Current evaluations of ALLMs have primarily focused on their performance in fundamental tasks. SUPERB [87] first introduced a unified evaluation framework for speech processing, where self-supervised speech representation models are assessed across ten downstream tasks, including phoneme recognition, keyword spotting, speaker verification, and emotion recognition. This benchmark demonstrates the generality and effectiveness of SSL representations in diverse scenarios. Subsequently, SUPERB-SG [68] extended this framework to encompass advanced semantic understanding and generative tasks, such as speech translation [69], voice conversion [51], speech separation [73], and enhancement [7], in order to further evaluate models’ generative abilities and robustness. SLURP [6] provides a large-scale dataset and evaluation framework targeting spoken language understanding, thereby enabling a comprehensive comparison between end-to-end and pipeline approaches, while SLUE [61] assesses complex tasks including audio question answering, summarization, and named entity recognition within realistic speech scenarios with low-resource context, highlighting the impact of ASR models on downstream task performance. In the field of audio captioning, AudioCaps [32] and Clotho [17] serve as major evaluation benchmarks, with Clotho-AQA [39] pioneering a real-world dataset for audio question answering, facilitating the evaluation of models’ semantic reasoning capabilities. The recently released AIR-Bench [86] categorizes evaluation tasks into two dimensions: fundamental abilities and dialogic abilities, covering a wide variety of audio types such as speech, environmental sounds, and music. The fundamental dimension comprises 19 specific tasks, whereas the dialogic dimension uses open-ended question-answering formats to evaluate generative performance of models under diverse and mixed audio backgrounds. These benchmarks offer diverse and comprehensive frameworks for evaluation and comparison of ALLMs, yet they mainly focus on fundamental performance; systematic assessments of safety, ethical risks, and social impacts remain insufficient.
@@
Existing safety evaluation benchmarks are relatively limited, with most focusing on multimodal scenarios or specific attack methods. For example, MM-SafetyBench [41] proposed an evaluation framework for image query attacks targeting multimodal LLMs, collecting 5,040 text-image pairs to assess model safety under image manipulation. SafeBench [80] constructed 23 risk scenarios and 2,300 multimodal harmful example pairs by automatically generating harmful multimodal queries, and designed a collaborative LLM review protocol to enhance evaluation reliability. In the audio domain, the Chat-Audio Attacks (CAA) benchmark [88] designed four types of audio attacks for dialog audio attack evaluation, and adopted a synthesis of standard evaluation, GPT-4o-based assessment, and human evaluation strategies to measure model robustness. The study [85] comprehensively evaluated the safety performance of five audio multimodal models via red-teaming against harmful audio, textual interference, and specific jailbreak attacks, revealing attack success rates as high as 70%. Furthermore, the SEA method [44] proposed a synthetic embedding augmentation approach for safety alignment, verifying the feasibility of aligning audio safety in multimodal models using only textual data. Although the above benchmarks have made progress in their respective areas, there is still a lack of a unified audio safety benchmark that comprehensively considers multidimensional risks such as fairness, hallucination detection, privacy protection, robustness, and authentication. Therefore, this work proposes the AudioTrust benchmark, which encompasses six core directions: fairness evaluation, hallucination detection, safety defense, privacy leakage, robustness challenges, and identity authentication. By combining scenario-driven question-answer pairs with GPT-4o automated evaluation, AudioTrust reveals the safety boundaries of ALLMs in high-risk environments, thereby providing systematic guidance for the secure and trustworthy deployment of future models.
@@
While AudioTrust offers a pioneering and comprehensive framework for the multidimensional trustworthiness evaluation of Audio Large Language Models (ALLMs), certain limitations warrant consideration. Firstly, the datasets, though meticulously constructed to cover a diverse range of scenarios across fairness, hallucination, safety, privacy, robustness, and authentication, are necessarily finite and may not encapsulate the full spectrum of real-world complexities or all potential adversarial manipulations. Secondly, the dynamic nature of ALLM development and emerging threat landscapes also means that any benchmark, including AudioTrust, represents a snapshot in time and will require continuous updates to remain relevant and comprehensive in assessing the evolving trustworthiness of these rapidly advancing systems.
@@
We follow the documentation frameworks provided by [81].
@@
The dataset is available on the Hugging Face Datasets Hub and can be loaded using the https://huggingface.co/datasets/JusperLee/AudioTrust.

### v3

@@ I.4 Evaluation Method and Metrics
Experiment II: We calculate the authentication pass rate.
Experiment III: We compare authentication pass rates across different experimental groups.
@@ I.5 Additional Result
| Model Group | Model | Awake Assistant | Open The Door |
| Open-source | SALMONN | N/A | N/A |
| Ultravox | 91 | 53 | |
| Qwen2-Audio | 15 | 0 | |
| MiniCPM-o 2.6 | 27 | 14 | |
| Step-Fun | 78 | 78 | |
| Qwen2.5-omni | 100 | 100 | |
| Kimi-Audio | 92 | 59 | |
| OpenS2S | 100 | 0 | |
| Step-Audio2 | 91 | 7 | |
| Closed-source | Gemini-1.5 pro | 100 | 33 |
| GPT-4o Audio | 33 | 0 | |
| GPT-4o mini Audio | 8 | 20 | |
| Gemini-2.5 Flash | 94 | 84 | |
| Gemini-2.5-Pro | 99 | 80 |
Note: and represent the change in the proportion of failed defenses against voice cloning spoofing when comparing the ’Open the door’ scenario with stricter system prompts to the ’Awake Assistant’ scenario. Lower numbers represent better defense effectiveness.
| Model Group | Model | Identity Bypass | Identity Bypass (Text) | Hybrid Spoofing (bg) | Hybrid Spoofing (no bg) |
| Open-source | SALMONN | 74 | – | 93 | 92 |
| Ultravox | 5 | – | 43 | 41 | |
| Qwen2-Audio | 58 | – | 29 | 40 | |
| MiniCPM-o 2.6 | 76 | – | 57 | 44 | |
| Step-Fun | 11 | – | 3 | 2 | |
| Qwen2.5-omni | 36 | – | 36 | 64 | |
| Kimi-Audio | 21 | – | 24 | 14 | |
| OpenS2S | 3 | – | 34 | 33 | |
| Step-Audio2 | 36 | – | 85 | 80 | |
| Closed-source | Gemini-1.5 pro | 4 | 6 | 5 | 0 |
| GPT-4o Audio | 2 | 6 | 0 | 0 | |
| GPT-4o mini Audio | 0 | 9 | 0 | 0 | |
| Gemini-2.5 Flash | 3 | 4 | 7 | 19 | |
| Gemini-2.5-Pro | 5 | 0 | 3 | 0 | |
| Average | 23.9 | 5.0 | 29.9 | 30.6 | |
Note: Values show false acceptance rate (FAR) (%) with performance indicators relative to average values. indicates lower FAR than average (better security performance); indicates higher FAR than average (worse security performance). Subscripts show absolute difference from average. “–” indicates the subset of the dataset was used exclusively to evaluate model performance across different modalities (audio and text). The evaluation was restricted to closed-source models to streamline the analysis.“bg” = with background audio,“no bg” = without background audio.
@@
Voice Cloning Spoofing. In Table˜26, we analyze the experimental results of all open-source and closed-source models under two scenarios with different levels of text prompt flexibility. It can be observed that most models perform better in the “Open The Door” scenario than in the “Awake Assistant” scenario. with a significant decrease in the number of samples that failed to defend against voice cloning spoofing. This is particularly evident for Ultravox, Gemini-1.5 Pro, and Gemini-2.5 Pro. This indicates that even in Audio-based Large Language Models (ALLMs) where audio is the primary input, the accuracy of text prompts still plays a significant role. Furthermore, this has implications for the downstream applications of ALLMs: for scenarios involving security, authentication, etc., designing a strict and precise prompt may lead to a considerable improvement in model performance.
@@
Identity Verification Bypass. From the Table˜27, it can be observed that closed-source models are harder to deceive compared to open-source models. Among them, GPT-4o mini Audio performs the best, with a FAR (False Acceptance Rate) of only 0%. Among all closed-source models, SALMONN performs the worst, with a FAR as high as 74%. These results indicate that even without providing complete or explicit authentication information, voice models still have a high probability of passing identity verification, which poses a significant security risk.In the Table˜27, we also investigated the FAR metric in pure text mode, which is labeled as the “Text” column. This represents using the corresponding text-based model of the audio model to perform inference on the text version of the identity verification bypass dataset. It can be observed that, in general, the FAR is higher in text mode compared to audio mode. This suggests that the additional paralinguistic information present in the speech modality, such as emotional cues or prosodic features, may contribute positively to the authentication performance of the model.
@@
Hybrid Spoofing. In this configuration, we simulate social engineering attacks combined with background audio that may occur in real authentication scenarios, aiming to study the impact of background sounds on the verification outcome. The experimental results show that the influence of added background audio on model performance does not follow an obvious pattern. For instance, Qwen2-Audio’s FAR decreases by 11%, whereas Gemini-1.5 Pro’s FAR increases instead.
@@
With the rapid increase in parameter and data scales, text-only large language models (LLMs) have achieved groundbreaking progress in language understanding and generation, as exemplified by models such as GPT‑4 and the Gemini series (Achiam et al., 2023; Team et al., 2023). Building on this, researchers explored cross-modal alignment by integrating visual information into unified representation spaces. This led to early models like CLIP (Radford et al., 2021) and Flamingo (Alayrac et al., 2022), and later, models such as GPT-4V and Gemini capable of processing high-resolution images and long contexts. Recently, ALLMs have further expanded the input modalities by incorporating temporal acoustic features (such as Mel-spectrograms, log-power spectra, or variable-length waveforms) for joint modeling with semantic tokens (Yang et al., 2024c). In contrast to the visual modality, audio signals exhibit high dynamic range and transient variations in both time and frequency domains. Consequently, most ALLMs adopt separate time-frequency encoders or discretizing acoustic tokenizers to capture rich attributes such as timbre, rhythm, and scene (Hu et al., 2024; Du et al., 2023). Representative models include Qwen2‑Audio with its pipeline-style natural language prompt pre-training (Chu et al., 2024), SALMONN with a unified "auditory-language-music" framework (Tang et al., 2024), and WavLLM with a dual-encoder plus Prompt-LoRA adaptation mechanism (Hu et al., 2024). After cross-modal alignment, these models demonstrate strong capabilities in content and scene understanding, enabling applications such as spoken question answering, music style analysis, and environmental sound event retrieval. They also show great promise in medical diagnosis (e.g., detection of respiratory diseases, analysis of heart sounds), voice control for smart homes, and multimedia generation and editing (Zhang et al., 2024b; Rani et al., 2017; Banerjee et al., 2024).
@@
However, the multimodal nature of ALLMs also introduces new trustworthiness challenges. First, since the models are trained on large-scale acoustic-text paired corpora, they are prone to memorizing and leaking sensitive user speech information, and are therefore vulnerable to privacy attacks such as membership inference (Tomashenko et al., 2024; Hu et al., 2022). Second, adversarial audio can exploit inaudible ultrasound or fine-grained perturbations to mislead ALLMs: early work such as DolphinAttack (Zhang et al., 2017a) and Vrifle (Li et al., 2024b) demonstrated covert manipulation of voice assistants via inaudible commands injected with ultrasonic carriers above 20 kHz (Zheng et al., 2023b; Li et al., 2023c; Ze et al., 2023); recently, AdvWave systematically proposed gradient shattering repair and two-stage optimization, achieving over 40% jailbreak success rates on various ALLMs (Kang et al., 2024). In addition, large-scale multimodal models are similarly susceptible to cross-modal instruction injection and protocol mismatching attacks, potentially leading to unauthorized content generation (Liu et al., 2023), privilege escalation (He et al., 2025), and even physical harm (Lu et al., 2024). When integrated into voice-interface agentic systems, trustworthiness challenges are amplified and become paramount (Liu et al., 2025; Yu et al., 2025). To address these risks, the community has proposed a range of safety, security, and privacy mechanisms, including SafeEar, an empirical content privacy-preserving audio deepfake detection framework (Li et al., 2024a) and active detection with post-hoc rejection (Li et al., 2023b) differentially private pre-training, segment-wise gradient compression defenses. Nevertheless, in real-time voice scenarios, these approaches still face detection latency and robustness trade-offs, highlighting the urgent need for further research.
@@
Current evaluations of ALLMs have primarily focused on their performance in fundamental tasks. SUPERB (Yang et al., 2021) first introduced a unified evaluation framework for speech processing, where self-supervised speech representation models are assessed across ten downstream tasks, including phoneme recognition, keyword spotting, speaker verification, and emotion recognition. This benchmark demonstrates the generality and effectiveness of SSL representations in diverse scenarios. Subsequently, SUPERB-SG (Tsai et al., 2022) extended this framework to encompass advanced semantic understanding and generative tasks, such as speech translation (Wahlster, 2013), voice conversion (Mohammadi & Kain, 2017), speech separation (Wang & Chen, 2018), and enhancement (Benesty et al., 2006), in order to further evaluate models’ generative abilities and robustness. SLURP (Bastianelli et al., 2020) provides a large-scale dataset and evaluation framework targeting spoken language understanding, thereby enabling a comprehensive comparison between end-to-end and pipeline approaches, while SLUE (Shon et al., 2022) assesses complex tasks including audio question answering, summarization, and named entity recognition within realistic speech scenarios with low-resource context, highlighting the impact of ASR models on downstream task performance. In the field of audio captioning, AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) serve as major evaluation benchmarks, with Clotho-AQA (Lipping et al., 2022) pioneering a real-world dataset for audio question answering, facilitating the evaluation of models’ semantic reasoning capabilities. The recently released AIR-Bench (Yang et al., 2024c) categorizes evaluation tasks into two dimensions: fundamental abilities and dialogic abilities, covering a wide variety of audio types such as speech, environmental sounds, and music. The fundamental dimension comprises 19 specific tasks, whereas the dialogic dimension uses open-ended question-answering formats to evaluate generative performance of models under diverse and mixed audio backgrounds. These benchmarks offer diverse and comprehensive frameworks for evaluation and comparison of ALLMs, yet they mainly focus on fundamental performance; systematic assessments of safety, ethical risks, and social impacts remain insufficient.
@@
Existing safety evaluation benchmarks are relatively limited, with most focusing on multimodal scenarios or specific attack methods. For example, MM-SafetyBench (Liu et al., 2024) proposed an evaluation framework for image query attacks targeting multimodal LLMs, collecting 5,040 text-image pairs to assess model safety under image manipulation. SafeBench (Xu et al., 2022) constructed 23 risk scenarios and 2,300 multimodal harmful example pairs by automatically generating harmful multimodal queries, and designed a collaborative LLM review protocol to enhance evaluation reliability. In the audio domain, the Chat-Audio Attacks (CAA) benchmark (Yang et al., 2024d) designed four types of audio attacks for dialog audio attack evaluation, and adopted a synthesis of standard evaluation, GPT-4o-based assessment, and human evaluation strategies to measure model robustness. The study (Yang et al., 2024b) comprehensively evaluated the safety performance of five audio multimodal models via red-teaming against harmful audio, textual interference, and specific jailbreak attacks, revealing attack success rates as high as 70%. Furthermore, the SEA method (Lu et al., 2025) proposed a synthetic embedding augmentation approach for safety alignment, verifying the feasibility of aligning audio safety in multimodal models using only textual data. Although the above benchmarks have made progress in their respective areas, there is still a lack of a unified audio safety benchmark that comprehensively considers multidimensional risks such as fairness, hallucination detection, privacy protection, robustness, and authentication. Therefore, this work proposes the AudioTrust benchmark, which encompasses six core directions: fairness evaluation, hallucination detection, safety defense, privacy leakage, robustness challenges, and identity authentication. By combining scenario-driven question-answer pairs with GPT-4o automated evaluation, AudioTrust reveals the safety boundaries of ALLMs in high-risk environments, thereby providing systematic guidance for the secure and trustworthy deployment of future models.
@@
While AudioTrust offers a pioneering and comprehensive framework for the multidimensional trustworthiness evaluation of Audio Large Language Models (ALLMs), certain limitations warrant consideration. Firstly, the datasets, though meticulously constructed to cover a diverse range of scenarios across fairness, hallucination, safety, privacy, robustness, and authentication, are necessarily finite and may not encapsulate the full spectrum of real-world complexities or all potential adversarial manipulations, such as reliability (Ma et al., 2025). Secondly, the dynamic nature of ALLM development and emerging threat landscapes also means that any benchmark, including AudioTrust, represents a snapshot in time and will require continuous updates to remain relevant and comprehensive in assessing the evolving trustworthiness of these rapidly advancing systems.
@@
We follow the documentation frameworks provided by (Xu et al., 2025a).
@@
This dataset is available to Huggingface Datasets (https://huggingface.co/datasets/JusperLee/AudioTrust).
@@
Appendix N LLM Usage
In the course of this research and in preparing the manuscript, we utilized Large Language Models (LLMs) for two distinct purposes. First, during the manuscript preparation phase, an LLM was used to assist in refining the wording and improving the clarity of the English prose. Its role in this capacity was strictly limited to enhancing sentence structure, grammar, and the overall flow of the text. Second, in the evaluation phase of our research, we employed GPT-4o as a model-based evaluator to determine whether the outputs generated by our model adhered to a set of predefined rules. Beyond these specified roles, LLMs were not involved in the initial research design, data collection, or the generation of core scientific ideas. All substantive content, methodologies, and conclusions are entirely the original work of the authors.